# Language Models Do Not Follow Occam’s Razor: A Benchmark for Inductive and Abductive Reasoning

### v1

@@ 3 Approach
Next, we generate observations so that LLMs have sufficient information to produce the ground truth hypotheses. The observations must be complex enough that the reasoning example is non-trivial, yet they should also ensure the reasoning is feasible.
@@
  • To generate an example where the task is to infer a property (e.g., All cats are cute), we first identify the type in the ontology with this property (e.g., cat), then we sample members of this type and subtypes (e.g., Amy is a cat; Sally is a ragdoll; Polly is a shorthair), and add observations where the members have this property (e.g., Amy is cute; Sally is cute; Polly is cute).
@@
  • To generate an example where the task is to infer a membership relation (e.g., Amy is a cat), we first identify the type in the ontology with this member (e.g., cat), then we sample properties of this type and supertypes (e.g., All cats are cute; All felines are not slow; All mammals are hairy), and add observations where this member have these properties (e.g., Amy is cute; Amy is not slow; Amy is hairy).
@@
  • To generate an example where the task is to infer a subtype relation (e.g., All cats are felines), we first identify the subtype in the ontology (e.g., cat), then we sample members of this type and its subtypes (e.g., Amy is a cat; Sally is a ragdoll; Polly is a shorthair), and add observations where these members belong to the supertype (e.g., Amy is a feline; Polly is a feline; Sally is a feline).
Translate theories and observations into natural language
@@
We use the same grammar as in PrOntoQA for translating logical forms into natural language. Properties are translated into the form: All cats are cute; subtype relations are randomly translated into one of three forms: Each ragdoll is a cat; Every ragdoll is a cat; All ragdolls are cats; membership relations are translated into the form Amy is a cat.
Hypothesis quality
@@
We define a metric to evaluate the quality of hypotheses based on the principles of parsimony, or Occam’s Razor (i.e., the priciple that, other things being equal, simpler theories are better;  Baker 2022). In this regard, we encourage one hypothesis to explain as many observations as possible and penalize unnecessary hypotheses. Formally, given a world model $\mathcal{T}$, a set of observations $\mathcal{O}$, their associated proof trees $\mathcal{P}$,candidate hypotheses $\mathcal{H}$, and ground truth hypotheses $\mathcal{H}^{*}$, for $h\in\mathcal{H}$, $n(h)$ denotes the number of times $h$ appears in the proof trees of $\mathcal{O}$ (i.e., how many times $h$ is used as a premise of a proof step). We use $|\mathcal{H}|$ to denote the cardinality of $\mathcal{H}$. The quality of $\mathcal{H}$, $q(\mathcal{H})$ is defined by  [^5] We write the equation as the ratio of two averages: the numerator is the average accuracy of the predicted hypothesis (i.e., the average number of times a hypothesis appears in any valid proof), and the denominator is the average number of times a hypothesis appears in the shortest (i.e., most parsimonious) proof.:
@@
According to the definition, the ground truth hypotheses $\mathcal{H}^{*}$ have a quality score of 1. Consider the example in Figure 1, here $\mathcal{H}^{*}$ is {Fae is a tiger, All mammals are hairy, All rodents are mammals} and $|\mathcal{H}^{*}|=3$. Each $h^{*}\in\mathcal{H}^{*}$ appears three times in the proof trees of $\mathcal{O}$, so $\sum_{h^{*}\in\mathcal{H}^{*}}n(h^{*})=9$. Consider candidate hypotheses $\mathcal{H}$ {Fae is a tiger, All mammals are hairy, All rats are mammals, All squirrels are mammals, All cats are hairy}. $|\mathcal{H}|=5$. Obviously $\mathcal{H}$ can explain $\mathcal{O}$. All rats are mammals appears twice, All squirrels are mammals appears once, and All cats are hairy appear once in the proof trees of $\mathcal{O}$, so $\sum_{h\in\mathcal{H}}n(h)=3+3+2+1+1=10$. In this case, $q(\mathcal{H})=\frac{3\times 10}{5\times 9}=0.67$.
Appendix C contains the complete algorithms for generating reasoning examples and evaluating hypotheses.
4 Results
@@ 4.1 Experimental setup
In each experiment configuration, we generate 100 independent and identically distributed reasoning examples. Each reasoning example is pure text, a user prompt with an incomplete world model and observations. We also specify the desired output format in the system prompt. To answer each reasoning example, LLMs must produce hypotheses to explain all observations. We evaluate hypotheses on three metrics: weak accuracy, strong accuracy, and quality. We compute 95% confidence intervals for both accuracy metrics using the Wilson interval (Wilson, 1927). We use the OpenAI API to evaluate GPT-4o and together.ai platform to evaluate Llama3-70B, Gemma3-27B, DeepSeek-V3, and DeepSeek-R1-Distill-Llama-70B.
In our experiments, we vary the following variables: (1) Different types of tasks (i.e., infer property, infer membership relation, or infer subtype relation), (2) Single hypothesis or multiple hypotheses, (3) The height of the ontology tree, (4) Whether the reasoning example contains in-context demonstrations and whether they are in-distribution or out-of-distribution.
@@ 4.2 Zero-shot results
To begin evaluating the inductive and abductive reasoning abilities of LLMs, we first investigate the simplest setting with zero-shot prompting.
@@ 4.2.1 Single hypothesis
We first investigate results for reasoning examples with one hypothesis for three tasks. This helps us to isolate different types of tasks and understand how LLMs perform in each case. The results are in Figure 2. In general, all models show high accuracy (both weak and strong) greater than 80% for all three tasks under a height-1 ontology tree. When the height of the ontology tree starts to increase, we also observe a general decreasing trend in all three metrics, which is not surprising. One interesting observation is that LLMs perform better in infer membership relation tasks than other tasks, especially for Gemma3-27B. For example, even under a height-4 ontology tree, Gemma3-27B can achieve near 50% weak accuracy, while the number drops to around 10% for infer subtype relation tasks. This may be because the weak correctness for infer membership relation tasks only requires searching for the concept that has some property, but do not need to understand the structure of the ontology tree.666e.g., given Amy is cute, LLMs only need to search for a concept which has the property cute to infer the membership relation of Amy. However, this is not the case for infer subtype relation and infer properties tasks. We observe a similar trend for infer subtype relation and infer properties tasks, as we can view the concept’s parent as a special property of that concept.
1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00QualityTask: infer membership relationTask: infer propertyTask: infer subtype relationGPT-4oLlama3-70BDeepSeek-V3Gemma3-27B
@@ 4.2.2 Multiple hypotheses
Next, we turn to multi-hypothesis scenarios because they are closer to real-life situations, and we are interested in understanding how LLMs perform in this setting. Each reasoning example can contain multiple hypotheses. To check the difficulty of each reasoning example, we sample the ground truth hypotheses such that the number of hypotheses grows linearly w.r.t. the height of the ontology tree. [^7] Each ground truth hypothesis has $\sim 3$ observations. So the number of observations also grows linearly w.r.t. the height of the ontology tree. Invariably, the number of axioms in the world model grows exponentially w.r.t. the height of the ontology tree. See Table 3 for statistics under multi-hypothesis scenarios and the top row of Figure 3 for results.
The height of the ontology tree # of world model axioms # of observations # of ground truth hypotheses 1 9.0 10.0 3.0 2 14.0 11.8 3.5 3 25.5 15.2 4.6 4 46.8 20.0 6.6
Table 3: Basic statistics of multi-hypothesis examples in InAbHyD . We list the average number of world model axioms, observations, and ground truth hypotheses w.r.t. the height of the ontology tree in the same set of reasoning examples we used in evaluation. Note that the number of the ground truth hypotheses and the number of observations grow linearly with the height of the ontology tree. In contrast, the number of world model axioms grows exponentially.
In general, the performance in the multi-hypothesis setting differed significantly from that in the single-hypothesis setting. For example, the accuracy of all models drops significantly when the height of the ontology tree is increased from 1 to 2, with accuracy going from above 80% to below 50%, except for GPT-4o. This is surprising because the average number of ground truth hypotheses only increases from $3.0$ to $3.5$. We attribute this to the coupling effect of multiple hypotheses to produce and the topology of the ontology tree. Reasoning examples under the ontology tree with height 3 or 4 achieve lower accuracies and quality scores for all models. We also observe a notable gap between weak accuracy and quality. For example, under the height-4 ontology tree, all models can achieve at least 20% weak accuracy, but struggle with extremely low quality and strong accuracy scores. We manually inspect some incorrect hypotheses and summarize the most common types of errors in Section 4.5.
@@ 4.3 Does in-context learning help?
Figure 3: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree with multiple hypotheses under zero-short (top), 8-shot in-distribution (middle), and 8-shot out-of-distribution (bottom). We list the absolute results of the 8-shot in-distribution and out-of-distribution in Appendix D.
LLMs have shown extraordinary generalization capabilities with few-shot examples coupled with CoT prompts (Brown et al., 2020). We are also interested in whether in-context learning can help with inductive and abductive reasoning. We add eight demonstrations to each reasoning example, with the demonstrations’ ground truth hypotheses and CoTs. The CoT consists of each observation’s step-by-step proof. We have two settings for in-context learning: (1) In-distribution demonstrations, where the demonstration has an ontology tree of the same height as the test question. (2) Out-of-distribution demonstrations, where each demonstration has a single hypothesis with a height-1 ontology tree. In both settings, we ensure demonstrations cover all types of tasks.
The in-distribution demonstrations are slightly more helpful than out-of-distribution demonstrations. This result is consistent with the hypothesis that in-context learning mimics Bayesian inference (Xie et al., 2022). We also observe that in-context learning with in-distribution demonstrations improves strong accuracy and quality in ontology trees with heights 3 and 4. We conjecture that LLMs can learn how to navigate the ontology hierarchy and produce high-quality hypotheses through in-context demonstrations. In contrast, we do not observe a significant performance improvement from out-of-distribution demonstrations.
4.4 Does RLVR help?
1234Height0.00.20.40.60.81.0Weak Accuracy1234Height0.00.20.40.60.81.0Strong Accuracy1234Height0.00.20.40.60.81.0QualityLlama3-70BDeepSeek-R1-Distill-Llama-70B
Figure 4: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree comparison results for Llama3-70B and DeepSeek-R1-Distill-Llama-70B with multiple hypotheses.
RLVR has demonstrated significant improvements in other reasoning benchmarks, and so we test whether this improvement generalizes to abductive and inductive reasoning. Specifically, we compare Llama3-70B and DeepSeek-R1-Distill-Llama-70B. In general, we observe an average performance improvement between 10% and 20%. To analyze why DeepSeek-R1-Distill-Llama-70B performs better than Llama3-70B, we also manually inspect the reasoning output of DeepSeek-R1-Distill-Llama-70B (i.e., tokens that are inside <think> and </think>). After producing hypotheses, DeepSeek-R1-Distill-Llama-70B will usually attempt to verify those hypotheses. This verification process is essentially a deductive reasoning process, and can catch some reasoning errors or fix low-quality hypotheses. A complete example is in Appendix E.
@@ 4.5 Analysis of common errors
To further understand how and why LLMs fail, we manually inspect mistaken and low-quality hypotheses. We summarize some of the most common types of errors here for further insights. We provide example responses in each category in Appendix E. (1) Wrong ontology direction.LLMs sometimes produce hypotheses in the wrong direction. For example, the ground truth hypothesis is All cats are mammals, while LLMs will produce the hypothesis All mammals are cats. (2) Ignore the ontology and produce unnecessary hypotheses.LLMs sometimes can produce unnecessary hypotheses because they ignore the ontology. For example, the ground truth hypothesis is All cats are cute, while LLMs produce two hypotheses: All cats are cute; All ragdolls are cute. (3) Fall back to trivial hypotheses.Technically, the observation is also a valid hypothesis, but it is considered low-quality because it can only explain itself. LLMs sometimes reuse observations as hypotheses. (4) Hallucinated entities.LLMs sometimes produce hypotheses with hallucinated entities (including concepts, properties, members). Usually, this happens when the entity appears in the in-context demonstrations, but LLMs use them in the test question. (5) Confuse members with concepts.We use fictional names for concepts and real-person names for members. However, LLMs sometimes do confuse these two. For example, the ground truth hypothesis is Amy is a wumpus, while LLMs produce Wumpus is Amy.
5 Conclusions and future work
In this study, we introduced a programmable and synthetic dataset, InAbHyD . Our dataset serves as an out-of-distribution evaluation of the inductive and abductive reasoning abilities of LLMs. We do not claim that success in our benchmark automatically implies success in real-world tasks, but rather, our benchmark serves as a checkpoint for developing models that are capable of robust real-world abductive and inductive reasoning.
@@ 5 Conclusions and future work
We systematically studied LLMs’ inductive and abductive reasoning capabilities. Our results suggest that current LLMs can’t generate high-quality hypotheses that match human-level intelligence. For example, suppose we use LLMs as a scientific discovery assistant. In that case, they may generate superficial hypotheses that can explain observations but fail to uncover more general hypotheses that reveal hidden structures or connections in science. Our results can serve to better inform practitioners of LLMs of their limitations, especially in applications that rely heavily on inductive and abductive reasoning. The dataset itself opens new avenues for future work, such as training or fine-tuning models to improve their inductive and abductive reasoning capabilities.
One limitation of this work is that first-order logic (FOL) limits the expressiveness of the world model and the reasoning example. We leave it as a future work to use higher-order logic (HOL) to generate more complex world models and reasoning examples, which would require translating from HOL into natural language and vice versa. The other crucial future direction is improving the quality of LLMs’ hypotheses. For that, LLMs may benefit from RLVR  (Lambert et al., 2024) that explicitly conveys principles of parsimony in inductive and abductive reasoning with an appropriate reward. Another interesting future direction will be to determine whether large language models fine-tuned on this synthetic dataset (using supervised fine-tuning or reinforcement learning from verifiable rewards) could improve their performance on real-world tasks involving inductive and abductive reasoning. If this is the case, this synthetic dataset is more valuable because we can essentially generate an infinite number of synthetic data points, without worrying about the bottleneck of sampling data from the real world.
Appendix A Detailed explanations of Figure 1
Here we list proof trees for each observation in Figure 1. Hypotheses are in green and world model axioms are in purple.
Fae is strong
[…]
Fae is not slow
$\forall x(\texttt{tiger}(x)\rightarrow\texttt{feline}(x))$     $\forall x(\texttt{feline}(x)\rightarrow\neg\texttt{slow}(x))$           $\forall x(\texttt{tiger}(x)\rightarrow\neg\texttt{slow}(x))$     $\texttt{tiger}(\texttt{Fae})$                             $\neg\texttt{slow}(\texttt{Fae})$
Fae is warm-blooded
$\forall x(\texttt{tiger}(x)\rightarrow\texttt{feline}(x))$     $\forall x(\texttt{feline}(x)\rightarrow\texttt{mammal}(x))$           $\forall x(\texttt{tiger}(x)\rightarrow\texttt{mammal}(x))$     $\forall x(\texttt{mammal}(x)\rightarrow\texttt{warm-blooded}(x))$                              $\forall x(\texttt{tiger}(x)\to\texttt{warm-blooded}(x))$     tiger(Fae)                                                     warm-blooded(Fae)
Sam is hairy
$\forall x(\texttt{mammal}(x)\to\texttt{hairy}(x))$     mammal(Sam)           hairy(Sam)
Alice is hairy
$\forall x(\texttt{cat}(x)\to\texttt{feline}(x))$     $\forall x(\texttt{feline}(x)\to\texttt{mammal}(x))$           $\forall x(\texttt{cat}(x)\to\texttt{mammal}(x))$     cat(Alice)                           mammal(Alice)     $\forall x(\texttt{mammal}(x)\to\texttt{hariy}(x))$                                                hairy(Alice)
Bob is hairy
$\forall x(\texttt{dog}(x)\to\texttt{canidae}(x))$     $\forall x(\texttt{canidae}\to\texttt{mammal}(x))$            $\forall x(\texttt{dog}(x)\to\texttt{mammal}(x))$     dog(Bob)                             mammal(Bob)     $\forall x(\texttt{mammal}(x)\to\texttt{hariy}(x))$                                                hairy(Bob)
Jack is a mammal
rat(Jack)     $\forall x(\texttt{rat}(x)\to\texttt{rodent}(x))$           rodent(Jack)      $\forall x(\texttt{rodent}(x)\to\texttt{mammal}(x))$                            mammal(Jack)
Noah is a mammal
rat(Noah)     $\forall x(\texttt{rat}(x)\to\texttt{rodent}(x))$           rodent(Noah)      $\forall x(\texttt{rodent}(x)\to\texttt{mammal}(x))$                            mammal(Noah)
Oliver is a mammal
squirrel(Oliver)     $\forall x(\texttt{squirrel}(x)\to\texttt{rodent}(x))$             rodent(Oliver)      $\forall x(\texttt{rodent}(x)\to\texttt{mammal}(x))$                                  mammal(Oliver)
Appendix B More examples

### v2

@@ 3.3 Observation generation
Next, we generate observations so that LLMs have sufficient information to produce the ground-truth hypotheses. The observations must be complex enough that the reasoning example is non-trivial, yet they should also ensure the problem is feasible.
@@
  • To generate an example where the task is to infer a property (e.g., All cats are cute), we first identify the type in the ontology with this property (e.g., cat), then we sample members of this type and subtypes (e.g., Amy is a cat; Sally is a ragdoll; Polly is a shorthair), and add observations where the members have this property (e.g., Amy is cute; Sally is cute;).
@@
  • To generate an example where the task is to infer a membership relation (e.g., Amy is a cat), we first identify the type in the ontology with this member (e.g., cat), then we sample properties of this type and supertypes (e.g., All cats are cute; All felines are not slow; All mammals are hairy), and add observations where this member have these properties (e.g., Amy is cute; Amy is not slow).
@@
  • To generate an example where the task is to infer a subtype relation (e.g., All cats are felines), we first identify the subtype in the ontology (e.g., cat), then we sample members of this type and its subtypes (e.g., Amy is a cat; Sally is a ragdoll; Polly is a shorthair), and add observations where these members belong to the supertype (e.g., Amy is a feline; Polly is a feline).
3.4 Naturalistic Question Generation
@@
We first use the same grammar as in PrOntoQA for converting logical forms into natural language using templates. Properties are converted into the form: All cats are cute; subtype relations are randomly converted into one of three forms: Each ragdoll is a cat; Every ragdoll is a cat; All ragdolls are cats; membership relations are converted into the form Amy is a cat. We then use GPT-4o to paraphrase it into more natural language. Details are provided in Appendix D.
3.5 Hypothesis quality
@@
We define a metric to evaluate the quality of hypotheses based on the principles of Occam’s Razor (i.e., the priciple that, other things being equal, simpler theories are better;  occ). In this regard, we encourage one hypothesis to explain as many observations as possible and penalize unnecessary hypotheses. Formally, given a world model $\mathcal{T}$, a set of observations $\mathcal{O}$, their associated proof trees $\mathcal{P}$,candidate hypotheses $\mathcal{H}$, and ground-truth hypotheses $\mathcal{H}^{*}$, for $h\in\mathcal{H}$, $n(h)$ denotes the number of times $h$ appears in the proof trees of $\mathcal{O}$ (i.e., how many times $h$ is used as a premise of a proof step). We first use $|\mathcal{H}|$ to denote the cardinality of $\mathcal{H}$. The quality of $\mathcal{H}$, $q(\mathcal{H})$ is defined by  [^4] We write the equation as the ratio of two averages: the numerator is the average accuracy of the predicted hypothesis (i.e., the average number of times a hypothesis appears in any valid proof), and the denominator is the average number of times a hypothesis appears in the shortest (i.e., most parsimonious) proof.:
@@
According to the definition, the ground-truth hypotheses $\mathcal{H}^{*}$ have a quality score of 1. Consider the example in Figure 1, here $\mathcal{H}^{*}$ is {Fae is a tiger, All mammals are hairy, All rodents are mammals} and $|\mathcal{H}^{*}|=3$. Each $h^{*}\in\mathcal{H}^{*}$ appears three times in the proof trees of $\mathcal{O}$, so $\sum_{h^{*}\in\mathcal{H}^{*}}n(h^{*})=9$. Consider candidate hypotheses $\mathcal{H}$ {Fae is a tiger, All mammals are hairy, All rats are mammals, All squirrels are mammals, All cats are hairy}. $|\mathcal{H}|=5$. Obviously $\mathcal{H}$ can explain $\mathcal{O}$. All rats are mammals appears twice, All squirrels are mammals appears once, and All cats are hairy appear once in the proof trees of $\mathcal{O}$, so $\sum_{h\in\mathcal{H}}n(h)=3+3+2+1+1=10$. In this case, $q(\mathcal{H})=\frac{3\times 10}{5\times 9}=0.67$.
We further motivate the hypothesis metric using Bayesian decision theory in Appendix E.1. We compare the metric to human evaluation of hypotheses in Appendix E.2. Appendix F contains the complete algorithms for generating reasoning examples and evaluating hypotheses.
4 Experiments
@@ 4.1 Experimental setup
In each experiment configuration, we generate 100 independent and identically distributed reasoning examples. Each reasoning example is pure text, a user prompt with an incomplete world model and observations. We also specify the desired output format in the system prompt. To answer each reasoning example, LLMs must produce hypotheses to explain all observations. We evaluate hypotheses on three metrics: weak accuracy, strong accuracy, and quality. We compute 95% confidence intervals for both accuracy metrics using the Wilson interval (wilson). We use the OpenAI API and together.ai platform.
In our experiments, we vary the following variables: (1) Different types of tasks (i.e., infer property, infer membership relation, or infer subtype relation), (2) Single hypothesis or multiple hypotheses, (3) The height of the ontology tree, (4) Whether the reasoning example contains in-context demonstrations and whether they are in-distribution (5) Different reasoning-augmented prompts such as buffer-of-thought (bot).
@@ 4.2 Zero-shot results
We first investigate the simplest setting with zero-shot prompting.
@@ 4.2.1 Single hypothesis
We first investigate results for reasoning examples with one hypothesis for three tasks. This helps us to isolate different types of tasks and understand how LLMs perform in each case. The results are in Figure 3. In general, all models show high accuracy (both weak and strong) greater than 80% for all three tasks under a height-1 ontology tree. When the height of the ontology tree starts to increase, we also observe a general decreasing trend in all three metrics, which is not surprising. One interesting observation is that LLMs perform better in infer membership relation tasks than other tasks, especially for Gemma3-27B. For example, even under a height-4 ontology tree, Gemma3-27B can achieve near 50% weak accuracy, while the number drops to around 10% for infer subtype relation tasks. This may be because the weak correctness for infer membership relation tasks only requires searching for the concept that has some property, but do not need to understand the structure of the ontology tree.555e.g., given Amy is cute, LLMs only need to search for a concept which has the property cute to infer the membership relation of Amy. However, this is not the case for infer subtype relation and infer properties tasks. We observe a similar trend for infer subtype relation and infer properties tasks, as we can view the concept’s parent as a special property.
12340.00.20.40.60.81.0Weak Accuracy12340.00.20.40.60.81.0Strong Accuracy12340.00.20.40.60.81.0Quality12340.00.20.40.60.81.0Weak Accuracy12340.00.20.40.60.81.0Strong Accuracy12340.00.20.40.60.81.0Quality1234Height0.00.20.40.60.81.0Weak Accuracy1234Height0.00.20.40.60.81.0Strong Accuracy1234Height0.00.20.40.60.81.0QualityTask: infer membership relationTask: infer propertyTask: infer subtype relationGPT-4oLlama3-70BDeepSeek-V3Gemma3-27B
@@ 4.2.2 Multiple hypotheses
Next, we turn to multi-hypothesis scenarios because they are closer to real-life situations, and we are interested in understanding how LLMs perform in this setting. Each reasoning example can contain multiple hypotheses. To check the difficulty of each reasoning example, we sample the ground-truth hypotheses such that the number of hypotheses grows linearly w.r.t. the height of the ontology tree. [^6] Each ground-truth hypothesis has $\sim 3$ observations. So the number of observations also grows linearly w.r.t. the height of the ontology tree. Invariably, the number of axioms in the world model grows exponentially w.r.t. the height of the ontology tree. See Table 3 in Appendix B for statistics under multi-hypothesis scenarios and the top row of Figure 4 for results.
Performance in the multi-hypothesis setting differed significantly from that in the single-hypothesis setting. For example, the accuracy of all models drops significantly when the height of the ontology tree is increased from 1 to 2, with accuracy going from above 80% to below 50%, except for GPT-4o. This is surprising because the average number of ground-truth hypotheses increases only from $3.0$ to $3.5$. We attribute this to the coupling effect of multiple hypotheses to produce and the topology of the ontology tree. Reasoning examples under the ontology tree with heights 3 or 4 achieve lower accuracy and quality scores across all models. We also observe a notable gap between weak accuracy and quality. For example, under the height-4 ontology tree, all models can achieve at least 20% weak accuracy but struggle with extremely low-quality and strong-accuracy scores. We manually inspect some incorrect hypotheses and summarize the most common types of errors in Section 4.5.
@@ 4.3 Does in-context learning help?
Figure 4: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree with multiple hypotheses under zero-shot (top), 8-shot in-distribution (middle), and 8-shot out-of-distribution (bottom). We list the absolute results of the 8-shot in-distribution and out-of-distribution in Appendix G.
LLMs have shown extraordinary generalization capabilities with few-shot examples coupled with CoT prompts (iclclassicla). We are also interested in whether in-context learning can help with inductive/abductive reasoning. We add $8$ demonstrations to each example, with the demonstrations’ ground-truth hypotheses and CoTs. The CoT consists of each observation’s step-by-step proof. We have two settings for in-context learning: (1) In-distribution demonstrations, where the demonstration has an ontology tree of the same height as the test question. (2) Out-of-distribution demonstrations, where each demonstration has a single hypothesis with a height-1 ontology tree. In both settings, we ensure demonstrations cover all types of tasks. We present our results in Figure 4.
The in-distribution demonstrations are slightly more helpful than out-of-distribution demonstrations. This result is consistent with the hypothesis that in-context learning mimics Bayesian inference (cotbay). We also observe that in-context learning with in-distribution demonstrations improves strong accuracy and quality in ontology trees with heights 3 and 4. We conjecture that LLMs can learn how to navigate the ontology hierarchy and produce high-quality hypotheses through in-context demonstrations. In contrast, we do not observe a significant performance improvement from out-of-distribution demonstrations.
4.4 Do reasoning models and reasoning-augmented prompts help?
Figure 5: Results for Llama3-70B and two large reasoning models gpt-5.4 and o3 with chain-of-thought prompting (top) and buffer-of-thought prompting (bottom).
Recent developments in large reasoning models, particularly those trained with RLVR, have led to substantial improvements across many evaluations. As shown in Figure 5, stronger models such as gpt5.4 and o3 consistently outperform LLaMA3-70B in both weak and strong accuracy and overall quality across different ontology tree heights. Moreover, reasoning-oriented prompting strategies further enhance performance: buffer-of-thought (BoT; bot) prompting generally yields higher scores than chain-of-thought prompting (CoT), especially in weak accuracy. However, despite these gains, a persistent gap remains between weak and strong accuracy across all models and settings—–for instance, weak accuracy often approaches near-perfect levels at lower heights, while strong accuracy remains notably lower and degrades more rapidly as height increases. This suggests that although recent advances and prompting techniques significantly improve reasoning capabilities, high-quality hypothesis generation remains an open challenge.
@@ 4.5 Analysis of common errors
To further understand how and why LLMs fail, we manually inspect mistaken and low-quality hypotheses. We summarize some of the most common types of errors here for further insights. We provide example responses in each category, quantitative analysis, and the details of the manual inspection process in Appendix H. (1) Wrong ontology direction.LLMs sometimes produce hypotheses in the wrong direction. For example, the ground-truth hypothesis is All cats are mammals, while LLMs will produce the hypothesis All mammals are cats. (2) Ignore the ontology and produce unnecessary hypotheses.LLMs sometimes can produce unnecessary hypotheses because they ignore the ontology. (3) Fall back to trivial hypotheses.Technically, the observation is also a valid hypothesis, but it is considered low-quality because it can only explain itself. LLMs sometimes reuse observations as hypotheses. (4) Hallucinated entities.LLMs sometimes produce hypotheses with hallucinated entities (including concepts, properties, members). Usually, this happens when the entity appears in the in-context demonstrations, but LLMs use them in the test question.
5 Conclusion
@@
In this study, we systematically examined the inductive/abductive reasoning capabilities of LLMs, especially whether they follow Occam’s Razor. Our results suggest that current LLMs can’t generate high-quality hypotheses that match human-level parsimony. Our results can serve to better inform practitioners of LLMs of their limitations, especially in applications that rely heavily on inductive/abductive reasoning. The proposed dataset opens new avenues for future work, such as training models to improve their reasoning capabilities.
Limitations
Our synthetic dataset serves as an out-of-distribution evaluation of the inductive and abductive reasoning abilities of LLMs. We do not claim that success in our benchmark automatically implies success in real-world tasks. Still, our benchmark serves as a checkpoint for developing models that are capable of robust real-world abductive and inductive reasoning.
Another limitation of this work is that first-order logic (FOL) restricts the expressiveness of the world model and the reasoning example. We leave it as future work to utilize higher-order logic (HOL) to generate more complex world models and reasoning examples.
Appendix A Preliminary background
A.1 Examples of different types of reasoning
For example, given “All humans are moral” and “Socrates is a human”, we can conclude that “Socrates is moral”. For example, if we see many cute cats, we may inductively reason that all cats are cute. For example, if we hear a bird sound outside, we may abductively reason that there is a bird. The latter two types of reasoning share a common high-level goal: to find explanations for observations. Furthermore, the composition of inductive and abductive reasoning is prevalent in real-world scenarios. For example, given (i) Amy has a skin rash and leg pain and has been diagnosed with a new, unknown virus. (ii) Sam has a skin rash and leg pain and has been diagnosed with the same new unknown virus. (iii) Tom also has a skin rash and leg pain. We can first induce the rule: The new unknown virus causes skin rash and leg pain, and then we can abduce that Tom is infected with this new virus.
A.2 Proof trees
A proof tree is a tree-structured representation of a logical derivation where each node corresponds to a statement, and each edge represents the application of an inference rule. The root of the tree is the final conclusion, while the leaves correspond to premises or axioms. Proof trees are commonly used in logic, theorem proving, and formal reasoning to make the structure of derivations explicit.
Formally, let $\Gamma$ denote a set of premises and $\varphi$ denote a conclusion. A proof tree demonstrates that $\Gamma\vdash\varphi$ by recursively applying inference rules until all leaf nodes correspond to elements of $\Gamma$ or axioms.
Example.
Consider the following premises:
$1.\quad$ $P\rightarrow Q$
$2.\quad$ $Q\rightarrow R$
$3.\quad$ $P$
We can derive $R$ using two applications of modus ponens. We first derive $Q$ given the premises $P$ and $P\rightarrow Q$. Next, we derive $R$ using the premises $Q$ and $Q\rightarrow R$. This proof can be represented as a proof tree, which is shown below.
$\frac{\frac{P\rightarrow Q\qquad P}{Q}\qquad Q\rightarrow R}{R}$
The tree therefore makes explicit the hierarchical structure of the reasoning process.
Appendix B Basic statistics of InAbHyD
We list basic statistics of InAbHyD in Table 1.
Ontology tree height # of world model axioms # of observations # of ground-truth hypotheses # of examples 1 9.0 10.0 3.0 500 2 14.0 11.8 3.5 500 3 25.5 15.2 4.6 500 4 46.8 20.0 6.6 500
Table 3: Basic statistics of multi-hypothesis examples in InAbHyD. We list the average number of world model axioms, observations, and ground-truth hypotheses w.r.t. the height of the ontology tree in the same set of reasoning examples we used in evaluation. Note that the number of the ground-truth hypotheses and the number of observations grow linearly with the height of the ontology tree. In contrast, the number of world model axioms grows exponentially.
Appendix C More examples of reasoning questions
In this section, we provide additional reasoning examples generated using our proposed framework. For each setting listed below, we provide an example consisting of a world model, list of observations, and ground-truth hypothesis/hypotheses.